Previous

Calendar

October

2009

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

27

JD270
Site Visit (SpB): Details , Photos (24)

Span cylinder at 820 psi. N2 at 1490 psi

mounted EOL csat-3 (sn 0674)

at 11:00 MST, started ops52/ops53; collecting eol csat3, accelerometer data no longer saved (after a few adjustments).

started running "fast_10.csi" in fast cr23x logger

moved EOL T/RH sensor up to 4.6m

A bit breezy in the morning, but calm and warm most of the day (very dry).

28

JD271

29

JD272

30

JD273

1

JD274
at approx 13:00 MST, power outage, data system goes down

2

JD275
Site Visit (SpB): Details , Photos (25)

Span cylinder at 750 psi. N2 at 1260 psi

at 9:30 MST, rebooted quacker (data system back up)

from 11:20-11:47 MST, Changed profiler (li-6251) dessicant

removed the KZ PAR-LITE sensor (PAR2.up) from the CNR logger

moved EOL T/RH sensor up to 8m

re-connected noahs snow depth logger and confirmed it was running

Very windy in the morning, but calm in the afternoon (very dry).

3

JD276

4

JD277
(Full "Harvest" Moon)

5

JD278
TRAM teardown week

6

JD279
Site Visit (SpB, LL, SO, KK, C, etc): Details , Photos (14)

Span cylinder at 700 psi. N2 at 1200 psi

helped move TRAM parts

removed extention cords and battery at West Robot

worked on grove2 (PAR teardown)

Very windy and cool in the morning, but calm in the afternoon.

7

JD280

8

JD281

9

JD282

10

JD283
Summer Jeep Rental (ends)


11

JD284

12

JD285

13

JD286
Yom Kippur

14

JD287

15

JD288

16

JD289
Site Visit (SpB, NT, DA, LY, CG, TD, NM, DG, PB, +): Details , Photos (18)

Span cylinder at 710 psi. N2 at 900 psi

moved UC radiation sensors back to mini-tower

re-connected UCB pressure sensor (ASCX15AN)

re-connected the mrs logger and resent the program (mrs_02.csi)

from 13:53-13:55, adjusted span/n2 flow rates for LI-6262

17

JD290

18

JD291

19

JD292

20

JD293

21

JD294

22

JD295

23

JD296

24

JD297

25

JD298

26

JD299
Site Visit (SpB, JS, T, A, P): Details

Span cylinder at 680 psi. N2 at 600 psi

Site visit with visitors from Taiwan.

3-4" of Fresh snow. Windy and Cold.

27

JD300
Site Visit (AT): Details

REA system removed from tower; also slow ozone sensor taken to NCAR for checking (AT).

28

JD301
Heavy Snow.

29

JD302
Heavy Snow.

30

JD303

31

JD304
Site Visit (SpB): Details , Photos (25)

Span cylinder at 520 psi. N2 at 550 psi

from 13:12-13:28 MST, Changed profiler (li-6251) dessicant

Sunny and windy (at times).

Prius Rental

Halloween